OpenStudy (anonymous):

what is the Find the area of a regular pentagon if a = 3 meters and s = 2 meters?

OpenStudy (anonymous):

sorry this is the question Find the area of a regular pentagon when a = 3 meters and s = 2 meters.

hartnn (hartnn):

whats the 's' ? a is the side right ?

OpenStudy (anonymous):

That's what I'm trying to figure out. ^^

OpenStudy (anonymous):

yes "a" is the side but I also dont know what is s?

hartnn (hartnn):

google gave me this formula \(\large A = (1/4) \sqrt{5(5+2\sqrt 5)}a^2\) just plug in a =3

OpenStudy (anonymous):

so A will be 15.484

OpenStudy (anonymous):

That's what I got.

hartnn (hartnn):

yep

OpenStudy (anonymous):

I'll post the give choices

OpenStudy (anonymous):

A) 15 square meters B) 18 square meters C) 24 square meters D) 30 square meters

hartnn (hartnn):

round it to 15 :)

OpenStudy (anonymous):

Does it say round the the nearest whole number?

hartnn (hartnn):

even if it doesn't say, we can.

OpenStudy (anonymous):

yes so its A but it doesn't say round it to the nearest whole #

OpenStudy (anonymous):

yes its the closest answer

OpenStudy (anonymous):

Yeah, it's the only answer choice that makes sense.

OpenStudy (anonymous):

@hartnn and @OrangeMaster thanks for your help :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

^^